# A Semantic-Aware and Multi-Guided Network for Infrared-Visible Image Fusion

### v2

@@ IV-D1 Setup
TABLE II: Detection comparisons of our method with eleven state-of-the-art methods on $M^{3}FD$ [4] dataset. Red indicates the best results, and blue shows the second-best values.
@@ IV-D2 Compared with SOTA methods
Tab. II and Fig. 8 display quantitative and qualitative results concerning object detection on $M^{3}FD$ dataset. SHIP[42] has the best detection performance on the truck. ReCoNet[37], SwinFusion[19] and CDDFuse[26] also show their strengths in detecting major objects in real street scenarios, especially for the cars and buses. Benefiting from the robust multi-guided architecture, the proposed model has the best detection performance in five classes, which demonstrates that generated fused images with specific characteristics can not only highlight targets difficult to detect but also preserve fundamental texture information.
@@ IV-E1 Setup
TABLE III: Segmentation comparisons of our method with nine state-of-the-art methods on $MFNet$[74] dataset. Red indicates the best results, and blue shows the second-best values.
@@ IV-E2 Compared with SOTA methods
The quantitative evaluation is shown in Tab. III. TarDal [4], SwinFusion [19] and CDDFuse [26] enable to efficiently classify unlabelled pixel regions. Besides, PMGI [18] has the highest segmentation accuracy. SwinFusion [19] and CDDFuse [26] have the highest value of mIoU on recognizing the car and person classes. Our proposed model has shown competitive results, especially in the bike and cone regions. By observing Fig. 9, Some methods, for example, DIDFuse [25], MURF [41], TarDal [4], U2Fusion [16] and PMGI [18] have difficulties in distinguishing the curves and SHIP [42] is hard to recognize car stops. SMFNet can better integrate the edge and contour information, and consequently makes semantic segmentation more accurate.
@@ IV-F Extension to Medical Image Fusion ($\mathbf{RQ3}$)
As an extension task to validate the generalizability of our methodology, we construct a dataset sourced from the Harvard Medical Website [^2] http://www.med.harvard.edu/AANLIB/home.html, including MRI-CT, MRI-PET and MRI-SPECT as training sets for the medical image fusion (MIF) task. There are 41 pairs of MRI-PET as testing sets. Since the hyperparameters are configured identically to those of the VIF task, there is no need for a validation set. The compared methods are CSMCA [75], NSST-PAPCNN [76], MATR [77], GeSeNet [78], MsgFusion [79], EMFusion [80] and IFCNN [17]. The fused medical images and objective evaluation results are listed in Fig. 10 and Tab. IV. The fused medical images compared in Fig. 10 demonstrate the superiority of the proposed model in persevering the structural information from MRI and the pseudo-color from PET. The objective evaluation results in Tab. IV further confirm that the proposed model can generate satisfactory fused images.
@@
TABLE V: Model complexity and running time on $MSRS$ dataset.
@@ IV-G Ablation Study ($\mathbf{RQ4},\mathbf{RQ5}$ and $\mathbf{RQ6}$)
In this subsection, we mainly validate the effectiveness of each module, feature decomposition and aggregation, and the reason for the two training stages. EN, SD, VIF and SSIM are employed as the objective metrics. Experiment results are shown in Tab. VI.
@@ IV-G1 Analysis on Modules
To validate the proposed CAI’s effectiveness, we replace it with the basic INN block (ID=AE1 in Tab. VI). The performance deteriorated as there was no cross-attention mechanism to enhance local semantic features. Note, different from the typical paradigm using cross attention to achieve cross-modal interaction, CAI is an implicit way to enhance high-frequency information. Additionally, we employed 4 stacked transformer blocks shown in Fig. 2 (d) without residual connections to capture low-frequency information in the encoder branch (ID=AE2). The results show that combining residual connections with CNNs improves long-range expression abilities. SMFNet without GNN (ID=AE3) leads to inadequate detailed feature extraction, which also affects performance across downstream tasks, as shown in Fig. 11. Lastly, in order to verify the proper fusion rules, we swapped two fusion layers (ID=AE4) to determine their optimal positions for the fusion stage.
TABLE VI: Ablation experiment results in the test set of RoadScene [16]. Red indicates the best results, and blue shows the second-best values.
ID Configurations EN SD VIF SSIM
AE1 from CAI to INN block 7.3257 49.5447 0.7139 0.9417
AE2 BFE w/o residual connections 7.3592 50.4884 0.7056 0.9291
AE3 w/o GNN 7.3692 50.9436 0.7027 0.9102
AE4 exchange detail and base fusion layer 7.3584 50.8266 0.7247 0.9482
AE5 w/o $\mathcal{L}_{semantic}^{I}$ 7.3579 50.7396 0.7297 0.9401
AE6 two-stage both with $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ 7.3765 51.4596 0.7248 0.9333
AE7 w/o two-stage training 7.365 50.4629 0.6906 0.9237
AE8 SMFNet 7.3909 51.7062 0.7328 0.9554
@@ IV-G2 Features Decomposition and Aggregation
In this part, we discuss the proposed semantic loss in training stage I (ID=AE5) and the correlation-driven GNN loss proposed in training stage II (ID=AE6). During the experiment, we found that two training stages both with the guidance of correlation loss function, made the fused images reduce information preservation and have low contrast. We hypothesize that GNN does not need explicit correlation restriction when executing the image reconstruction task. And in the fusion stage, it’s necessary to increase correlations between nodes $\mathbf{C}$ and $\mathbf{D}$ (explained in Sec. III-B4) as detail features. Gram matrix brings salient targets and maintains many texture details. $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ in the second stage further enhances cross-modality interaction and balance between high-frequency and low-frequency information.
@@ IV-G3 Two Training Stages
To validate the importance of two training stages, we directly train the encoder, decoder and fusion layers together (ID=AE7) with 120 epochs. The results are unsatisfactory. Pre-trained encoder and decoder show more powerful feature extraction abilities and promote training in the second stage to generate better performance on qualitative and quantitative results.
IV-H Efficiency Comparison
Tab. V lists the computational costs and execution time of the twelve image fusion methods on the 361 image pairs of the MSRS test set with the size of $640\times 480$. MURF, U2Fusion and PMGI run on TensorFlow platform with CPU. Other methods run on Pytorch with GPU acceleration. By numerical comparison, U2Fusion and PMGI have the minimal computational load. Additionally, The model of ReCoNet only has 8K parameters. DenseFuse and DIDFuse spend an average of 0.01 seconds generating per fused image. The proposed SMFNet spends less generating time than SwinFusion and has fewer parameters than DeFusion.
@@ V CONCLUSION
In this paper, a semantic-aware multi-guided network for infrared-visible image fusion was proposed. Graph Reasoning module models and reasons about high-level relations between two modalities and extract low-level detail features. Cross Attention and Invertible Block (CAI) solves information loss during the decomposing process to extract high-frequency features. Additionally, a base feature extraction module equipped with Multi-Dconv Head Transposed Attention (MDTA) and Local-enhanced Feed Forward (LeFF) network is primarily utilized to capture global features with long-range dependencies. Two novel loss functions are proposed as the restriction in two training stages. Experiments demonstrate the fusion effect of our proposed SMFNet and the accuracy of downstream pattern recognition tasks can also be improved. In further works, we will explore multi-task learning (MTL) including image reconstruction, fusion tasks and scene parsing, considering novel joint-optimization among these tasks soon.

### v3

@@ IV-D1 Setup
TABLE III: Detection comparisons of our method with eleven state-of-the-art methods on $M^{3}FD$ [4] dataset. Red indicates the best results, and blue shows the second-best values.
@@ IV-D2 Compared with SOTA methods
Tab. III and Fig. 8 display quantitative and qualitative results concerning object detection on $M^{3}FD$ dataset. SHIP[38] has the best detection performance on the truck. ReCoNet[36], SwinFusion[18] and CDDFuse[22] also show their strengths in detecting major objects in real street scenarios, especially for the cars and buses. Benefiting from the robust multi-guided architecture, the proposed model has the best detection performance in five classes, which demonstrates that generated fused images with specific characteristics can not only highlight targets difficult to detect but also preserve fundamental texture information.
@@ IV-E1 Setup
TABLE IV: Segmentation comparisons of our method with nine state-of-the-art methods on $MFNet$[72] dataset. Red indicates the best results, and blue shows the second-best values.
@@ IV-E2 Compared with SOTA methods
The quantitative evaluation is shown in Tab. IV. TarDal [4], SwinFusion [18] and CDDFuse [22] enable to efficiently classify unlabelled pixel regions. Besides, PMGI [17] has the highest segmentation accuracy. SwinFusion [18] and CDDFuse [22] have the highest value of mIoU on recognizing the car and person classes. Our proposed model has shown competitive results, especially in the bike and cone regions. By observing Fig. 9, Some methods, for example, DIDFuse [21], MURF [37], TarDal [4], U2Fusion [15] and PMGI [17] have difficulties in distinguishing the curves and SHIP [38] is hard to recognize car stops. SMFNet can better integrate the edge and contour information, and consequently makes semantic segmentation more accurate.
@@ IV-F Extension to Medical Image Fusion ($\mathbf{RQ3}$)
As an extension task to validate the generalizability of our methodology, we construct a dataset sourced from the Harvard Medical Website [^2] http://www.med.harvard.edu/AANLIB/home.html, including MRI-CT, MRI-PET and MRI-SPECT as training sets for the medical image fusion (MIF) task. There are 41 pairs of MRI-PET as testing sets. Since the hyperparameters are configured identically to those of the VIF task, there is no need for a validation set. The compared methods are CSMCA [73], NSST-PAPCNN [74], MATR [75], GeSeNet [76], MsgFusion [77], EMFusion [78] and IFCNN [16]. The fused medical images and objective evaluation results are listed in Fig. 10 and Tab. V. The fused medical images compared in Fig. 10 demonstrate the superiority of the proposed model in persevering the structural information from MRI and the pseudo-color from PET. The objective evaluation results in Tab. V further confirm that the proposed model can generate satisfactory fused images.
@@
TABLE VI: Model complexity and running time on $MSRS$ dataset.
@@ IV-G Ablation Study ($\mathbf{RQ4},\mathbf{RQ5}$ and $\mathbf{RQ6}$)
In this subsection, we mainly validate the effectiveness of each module, feature decomposition and aggregation, and the reason for the two training stages. EN, SD, VIF and SSIM are employed as the objective metrics. Experiment results are shown in Tab. VII.
@@ IV-G1 Analysis on Modules
To validate the proposed CAI’s effectiveness, we replace it with the basic INN block (ID=AE1 in Tab. VII). The performance deteriorated as there was no cross-attention mechanism to enhance local semantic features. Note, different from the typical paradigm using cross attention to achieve cross-modal interaction, CAI is an implicit way to enhance high-frequency information. Additionally, we employed 4 stacked transformer blocks shown in Fig. 2 (d) without residual connections to capture low-frequency information in the encoder branch (ID=AE2). The results show that combining residual connections with CNNs improves long-range expression abilities. SMFNet without GNN (ID=AE3) leads to inadequate detailed feature extraction, which also affects performance across downstream tasks, as shown in Fig. 11. Besides, we changed the addition $\oplus$ operation at the end of the encoder into the concatenation C operation (ID=AE4). The C operation increases computational cost and adds complexity of handling high-dimensional features. Lastly, in order to verify the proper fusion rules, we swapped two fusion layers (ID=AE5) to determine their optimal positions for the fusion stage.
TABLE VII: Ablation experiment results in the test set of RoadScene [15]. Red indicates the best results, and blue shows the second-best values.
ID Configurations EN SD VIF SSIM mAP@0.5
AE1 from CAI to INN block 7.3257 49.5447 0.7139 1.3287 0.818
AE2 BFE w/o residual connections 7.3592 50.4884 0.7056 1.3023 0.82
AE3 w/o GNN 7.3692 50.9436 0.7027 1.2896 0.810
AE4 from addition to concatenation 7.3773 50.7761 0.7893 1.2982 0.821
AE5 exchange detail and base fusion layer 7.3584 50.8266 0.7247 1.3265 0.818
AE6 w/o $\mathcal{L}_{semantic}^{I}$ 7.3579 50.7396 0.7297 1.325 0.829
AE7 two-stage w/o $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ 7.338 49.8241 0.7329 1.33 0.825
AE8 two-stage both with $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ 7.3765 51.4596 0.7248 1.3165 0.824
AE9 w/o two-stage training 7.365 50.4629 0.6906 1.3102 0.819
AE10 $\alpha_{1},\alpha_{3}=1$ 7.356 50.4861 0.7188 1.3303 0.83
AE11 $\alpha_{1},\alpha_{3}=5$ 7.3788 50.8299 0.7126 1.3168 0.825
AE12 $\alpha_{1},\alpha_{3}=8$ 7.3732 51.3246 0.7265 1.333 0.829
AE13 $\alpha_{1},\alpha_{3}=10$ 7.3791 50.9621 0.6999 1.2858 0.816
AE14 SMFNet 7.3909 51.7062 0.7328 1.33 0.835
Figure 12: Feature decomposition results on $MSRS$ dataset. The three rows correspond to “00706N”, “01603D” and “01110N” (from top to bottom).
@@ IV-G2 Features Decomposition and Aggregation
In this part, we discuss the proposed semantic loss in training stage I (ID=AE6) and the correlation-driven GNN loss proposed in training stage II (ID=AE7 and AE8). During the experiment, we found that the absence of $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ to guide the fusion of GNN features in the second stage hinders the fusion performance. As a comparison, two training stages both with the guidance of correlation loss function, made the fused images reduce information preservation and have low contrast. We hypothesize that GNN does not need explicit correlation restriction when executing the image reconstruction task[79]. And in the fusion stage, it’s necessary to increase correlations between nodes $\mathbf{C}$ and $\mathbf{D}$ (explained in Sec. III-B4) as detail features. Gram matrix brings salient targets and maintains abundant texture details. $CC(\Phi_{G}^{V},\Phi_{G}^{I})$ in the second stage further enhances cross-modality interaction and balance between high-frequency and low-frequency information. From Fig. 12, we provide the feature maps extracted from the three-branches encoder. Base features $\Phi_{B}^{I}$ and $\Phi_{B}^{V}$ contain gradual intensity changes with fewer sharp edges. Detail features $\Phi_{D}^{I}$ and $\Phi_{D}^{V}$ capture abrupt intensity changes and emphasize local contrast. $\Phi_{G}^{I}$ and $\Phi_{G}^{V}$ extracted from GR module highlight the edge and boundary of the objects and are also as the complementary of detail features.
@@ IV-G3 Two Training Stages
To validate the importance of two training stages, we directly train the encoder, decoder and fusion layers together (ID=AE9) with 120 epochs. The results are unsatisfactory. Pre-trained encoder and decoder show more powerful feature extraction abilities and promote training in the second stage to generate better performance on qualitative and quantitative results. To validate the weighting of decomposition loss, we set $\alpha_{1},\alpha_{3}=1,5,8,10$, respectively (from ID=AE10 to ID=AE13). The settings we chose $\alpha_{1},\alpha_{3}=2$ is optimal. The $\mathcal{L}_{decomp}^{I}$ and $\mathcal{L}_{decomp}^{II}$ play key roles in guiding the network to obtain decomposition features for the reconstruction and fusion procedure.
TABLE VIII: Quantitative evaluation results on $MSRS$ dataset compared with big model based methods, where red indicates the best results, and blue shows the second-best values.
Method EN SD SF MI VIF Qabf AG SSIM GFlops Params Time
Text-IF[80] 6.7281 44.5889 11.8784 2.8202 1.0311 0.6747 5.4684 1.3913 1519 89014 0.21
Text-DiFuse[81] 7.1435 54.2258 11.406 1.9698 0.7306 0.4406 5.212 0.9602 8604 119398 15
SMFNet 6.7021 42.9562 11.6976 3.1131 1.0448 0.7097 5.4765 1.3905 1462 4504 0.32
IV-H Discussion
IV-H1 Efficiency Comparison
Tab. VI lists the computational costs and execution time of the twelve image fusion methods on the 361 image pairs of the MSRS test set with the size of $640\times 480$. MURF, U2Fusion and PMGI run on TensorFlow platform with CPU. Other methods run on Pytorch with GPU acceleration. By numerical comparison, U2Fusion and PMGI have the minimal computational load. Additionally, The model of ReCoNet only has 8K parameters. DenseFuse and DIDFuse spend an average of 0.01 seconds generating per fused image. The proposed SMFNet spends less generating time than SwinFusion and has fewer parameters than DeFusion. In Tab. VIII, Text-IF [80] and Text-DiFuse[81] have the largest model parameters and computational cost owing to the help of big model, e.g., CLIP[80] and diffusion model[32].
IV-H2 More insights on our model
SMFNet exhibits stable performance on infrared-visible image fusion task and the fused images can also be applied into downstream tasks. In this paper, we decomposed features into high-frequency and low-frequency features by distinct architecture (i.e., CNN, Transformer and GNN). The correlation coefficient constrains features of the same type from two modalities to enhance detail preservation and saliency information in a multi-guided manner. Recent works proposed an overlooked issue: the over-reliance on visual features in the realm of image fusion[82]. Some pioneers[82], [83], [80], [81] extracted textual semantic information and guided the extraction and fusion procedures at the vision level. In Tab. VIII, the big model guided methods quantitative results on MSRS dataset are listed as a fair comparison. Notably, a complex network with high computational cost is challenging to be deployed on small-footprint devices, e.g., mobile phones. The utilization of text features and the computational overhead are our concerns to focus on for further improvements.
@@ V CONCLUSION
In this paper, a semantic-aware multi-guided network for infrared-visible image fusion was proposed. Graph Reasoning module models and reasons about high-level relations between two modalities and extract low-level detail features. Cross Attention and Invertible Block (CAI) solves information loss during the decomposing process to extract high-frequency features. Additionally, a base feature extraction module is primarily utilized to capture global features with long-range dependencies. Two novel loss functions are proposed as the restriction in two training stages. Experiments demonstrate the fusion effect of our proposed SMFNet and the accuracy of downstream pattern recognition tasks can also be improved. In further works, we will explore multi-task learning (MTL) including image reconstruction, fusion tasks and scene parsing, considering novel joint-optimization among these tasks soon.